Wavenumbers of yellow radiation with wavelength 5800A°
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
Wave number = 1/Wavelength = 1/(5800 Å)
= 1724137.93 m^-1

Frequency = c/Wavelength
= (3×10^8)/(5800)
= 5.168×10^14 Hz